OpenStudy (anonymous):

i have a question in one sided limits ? anyone to help . i can post the question

OpenStudy (anonymous):

anyone pl

OpenStudy (anonymous):

the first question

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

whats your doubt ? you can just plug in x=2 in left and right side functions and if the value is same, the limit exist

OpenStudy (anonymous):

is it that all about ?

OpenStudy (anonymous):

arent we supposed to do like this ?

hartnn (hartnn):

f is continuous if left side limit = right side limit

OpenStudy (anonymous):

thank you ! so just plug in the value ? thats it ?

hartnn (hartnn):

yes, you just need to select the proper function like for \(x \to 2^-\) since x<2, you will select "x-4" and then plug in x=2 (you can directly substitute as you are not getting any indeterminate form) for \(x \to 2^+\), you'll select x^2-6 and plug in x=2

hartnn (hartnn):

we need to plug in x=2, as we are considering the continuity at x=2 apart from x=2, the function is continuous everywhere, because x-4 and x^2-6 are continuous everywhere

OpenStudy (anonymous):

Thank you so much @hartnn i got it..

hartnn (hartnn):

welcome ^_^ you seem to be new here, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

OpenStudy (anonymous):

yes i just logged it,,, Thank you for the help and many more questions waiting,,!!

hartnn (hartnn):

bring it on! :)

OpenStudy (anonymous):

question 20 pl

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

1/x is infinity at x=0+ and -infinity at x=0- so its not continuous at x=0 only

OpenStudy (anonymous):

Thank you !!

OpenStudy (anonymous):

so apart from x=0, its continous at all values ?

OpenStudy (anonymous):

in the question , they never asked about x=0 right ?

hartnn (hartnn):

yes. oh yeah, then f is continuous everywhere

OpenStudy (anonymous):

What will be the proper answer to this ? when i get to write at exams

OpenStudy (anonymous):

i mean the step by step procedure @hartnn

OpenStudy (precal):

You just state that the domain is all real numbers except 0